OpenStudy (anonymous):

Find the slope of the line whose equation is px - qy = r + 1

OpenStudy (anonymous):

If you can get it into the form: y = (something)x + (somethingelse) your slope will be whatever "something" is because it will be in slope intercept form: y = mx + b

OpenStudy (mathstudent55):

Solve the equation for y. Then, then coefficient of x is the slope.

OpenStudy (anonymous):

px - qy = r + 1 => qy = px - (r+1) => y = px/q - (r+1)/q = slope*x + c => Slope = p/q.

OpenStudy (anonymous):

thanks!